A car is traveling at a constant speed of 27.3 m/s on a highway. At the instant this car passes an entrance ramp, a second car enters the highway from the ramp. The second car starts from rest and has a constant acceleration. What acceleration must it maintain, so that the two cars meet for the first time at the next exit, which is 2.06 km away?


1
Expert's answer
2021-10-08T08:11:03-0400
"vt=0.5at^2=d\\
a=\frac{2d}{t^2}\\
a=\frac{2v^2}{d}\\a=\frac{2(27.3)^2}{2060}=0.724\frac{m}{s^2}"